How do I do problems like this?
Solve.
√y+16=1
Please help me?


Answer by stanbon(75887)

√y+16=1
----------
sqrt(y+16) = 1
Square both sides to get:
x+16= 1
x=-15
---------
Check your answer:
sqrt(-15+16)=1
sqrt(1)= 1
1=1
==============
Cheers,
Stan H>